Calculate the mole fraction of phosphoric acid (H3PO4) in a 25.4% (by mass) aqueous solution. (Assume 750 mL of solution.) What is the molarity of the solution? What is the molality? (At 20 ° C, the density of phosphoric acid is 1.1462 g/mL and the density of water is 0.99823 g/mL.)
Answer #1

Answer let we have 100 g aquous solution. then then 100g 25.4g H₂Pou Moleculare = (3+1) +31 +(16x4) = 984 74.69 H₂O MolecularMolarity of He pou = moles of H₃PO4 x 1000 volume of solution (in ml) 0.259 x1000 750 = Molality of H₂PO = 259 0.345 M 750 mounit of molarity = mol/litre

Unit of molality = mol/kg

In question mass percentage is given so I assumed we have total 100 gram solution and the calculations done accordingly.